OpenStudy (kpop_fan):

Help please! A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate cookie, 3 almond cookies, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking 1 almond cookie followed by another almond cookie, and show the equation used.

OpenStudy (crabbyoldgamer):

Probability is desired/total. In the first draw, there are 3 almond cookies (desired) and 6 total cookies. The probability of an almond cookie is 3/6 = 1/2, In the second draw there are 2 almond cookies and 5 total cookies. The probability of an almond cookie is 2/5. To find the probability of almond cookie then almond cookie, multiply those two probabilities: 1/2 * 2/5 = 2/10 = 1/5 = 20%